Two ice skaters stand together hands to hands and "push off" so that they travel in exactly opposite directions. If the boy's weight is 735 N and the girl's weight is 490 N, what is the girl's velocity if the boy's velocity is 0.50 m/s to the left?
Answer #1

Given that, Weight of the boy = 735 N

Hence, His mass will be, m1 = (weight) /(g) = (735 N) / (9.8 m/s2) = 75 kg

Also, her velocity, v1 = - 0.50 m/s

(I'm taking left as negative and right as positive)

Wight of the girl = 490 N

Hence, Her mass will be, m2 = (weight) /(g) = (490 N) / (9.8 m/s2) = 50 kg

Apply law of conservation of momentum.

Total initial momentum = total final momentum

0 = m1v1 + m2v2

Where, v2is the final speed of the girl.

0 = (75 * -0.50) + (50) *v2

v2 = (37.5) / (50)

v2 = 0.75 m/s

  • That is, speed of the girl is 0.75 m/s to the right.

(in case of anything wrong/have any doubts, please reach out to me via comments. I will help you)